OpenStudy (anonymous):

Lisa drives to post-office at 40 miles per hour and returns though the same route at 50 miles per hour. If the return trip takes 15 minutes shorter, how long is the trip to the post office?

OpenStudy (anonymous):

We know the equation v = d /t Let d be the distance she travels, and is the same on the way there and back. For the first trip, 1) 40 = d / t Let t be the time she needs to take to get to the post office And for the second trip, 2) 50 = d / t -15 Isolating for d, 1) d = 40t 2) 50(t-15) = d Equating 1) and 2): We get 40t = 50(t-15) Solving for t gets you the time it takes for her to get to the post office

OpenStudy (dumbcow):

Just a note: since speed is in mph, time must be in units of hours 40t = 50(t-.25)